# A Prospect-Theoretic Policy Gradient Framework for Behaviorally Nuanced Reinforcement Learning

### v3

@@ 5 Simulations
Setting: We train a softmax policy using CPT-PG, vPG and ERS-PG. Vanilla PG optimizes for the standard expected return (identity for $u$ and $w$), CPT-PG using a KT model with S-shaped utility using parameters ($\alpha=0.6$, $\lambda=2.5$) and an S-shaped probability function $w$ under-weighting the probability 1/2. ERS-PG is run using an exponential (concave) risk sensitive utility $u(x)=\eta^{-1}(1-\exp(-\eta x))$ with $\eta=0.5\,.$ We use the exact same parameters for both environments.
Results and interpretation: As shown in Fig. 1 (right), in the Gain Bandit vPG selects the risky arm (higher expected return), whereas both CPT-PG and ERS-PG choose the safe arm—reflecting human risk aversion over gains. In the Loss Bandit, only CPT-PG “flips” to the risky arm, capturing human risk-seeking over losses. Crucially, CPT-PG does this with the same S-shaped utility and probability-weighting parameters in both settings, whereas the concave ERS-PG objective remains risk-averse throughout. This exact “safe in gains, risky in losses” pattern—known as the reflection effect—is the hallmark of Prospect Theory (Kahneman and Tversky, 1979; Tversky and Kahneman, 1992) from which our simple example is inspired. Neither expected-utility nor a single-parameter risk-sensitive criterion can reproduce both behaviors simultaneously, but CPT-RL can, using a fixed, psychologically-grounded model. While risk-sensitive RL can, with appropriate parameter tuning, replicate some aspects of human decision-making, it often lacks the asymmetric treatment of gains and losses and the probabilistic distortion that characterize human behavior. Overall our main message is that CPT-RL offers more flexibility in modeling. Note also that CPT captures several risk measures as particular cases using discontinuous weights (see App. G.3G.4).
(b) Robustness to state space size. We compare our PG algorithm with the zeroth-order (CPT-SPSA-G) of L.A. et al. (2016) on MDPs with increasing size $n\times n$. Fig. 2 shows that our CPT-PG algorithm scales better to larger grid sizes than CPT-SPSA-G as expected due to its use of (first-order) gradient information. See App. K.5 for details.
@@
(c) Application to finance. The goal is to train RL trading agents using our PG algorithm in the CPT-RL setting using a gym trading environment and data from the Bitcoin USD market. See App. K.8 for more details. We test several utility and probability weighting functions including a risk averse exponential of the form $x\mapsto\frac{1}{\beta}(1-\exp(-\beta x))$ with different values of $\beta$ as well as the KT (Kahneman and Tversky) function with different values of the reference point $x_{0}$ to illustrate its influence. In Figure 3, we make three observations. First, the reference point shifts the values of the achieved CPT returns: The smaller the reference point, the larger are the returns (Fig. 3, left). This is because only values larger than the reference point are perceived as positive returns. This illustrates how the subjective perception of the agent of the returns is taken into account by the model. Second, different values of $\beta$ lead to different trajectories overall which can translate to different levels of risk aversion. In particular, the curves do not match the identity utility case in the first episodes and show more or less risk taken towards optimizing the CPT returns. Third, the exponent $\alpha$ in the utility distorts the function and shifts the returns significantly (Fig. 3, right). Lower values of $\alpha$ lead to higher returns in this setting. This parameter $\alpha$ provides a degree of freedom to model the behavior of the agent as per their perception of the returns. Different values of $\alpha$ modify the curvature of the utility function (w.r.t. the reference point $x_{0}=0$ here) which is concave for gains and convex for losses.
@@
Additional simulations. We provide more simulations demonstrating the applicability of our CPT-PG algorithm in different settings, including illustrations of our theoretical results (App. K.3-K.4) and applications to a traffic control example over a grid (App. K.6), an electricity management example (App. K.7) and a control application (App. K.9).
@@ 6 Related Work
Prospect Theory and its sibling, CPT (Kahneman and Tversky, 1979; Tversky and Kahneman, 1992; Barberis, 2013), were first integrated with RL by L.A. et al. (2016). Since then, only a few studies have explored the CPT-RL framework (Borkar and Chandak, 2021; Ramasubramanian et al., 2021; Ethayarajh et al., 2024). Notably, Borkar and Chandak (2021) proposed a Q-learning algorithm for CPT-based policy optimization, while Ramasubramanian et al. (2021) developed value-based methods estimating CPT values using dynamic programming. Their approach optimizes a sum of CPT-transformed period costs, making it amenable to dynamic programming (see remark 1 therein). In contrast, our CPT formulation is different: we maximize the CPT value of the return of a policy CPT-PO. This objective lacks an additive structure, hence does not satisfy a Bellman equation, rendering dynamic programming approaches inapplicable. Additionally, prior value-based methods are limited to finite state-action spaces, whereas our PG algorithm is also suitable for continuous state action settings, as shown in our experiments. More recently, Ethayarajh et al. (2024) incorporated CPT (without probability distortion) for fine-tuning large language models with human feedback. Our work complements prior CPT-RL studies (L.A. et al., 2016; Jie et al., 2018) that rely on zeroth-order SPSA methods (Spall, 1992). Instead, we introduce a PG algorithm that leverages first-order information, exploiting the structure of CPT values applied to cumulative rewards (see Section 3 for further comparison). Unlike existing PG approaches in risk-sensitive RL, our method explicitly accounts for probability distortion and S-shaped utility transformations, key aspects of CPT. For a broader discussion on CPT-RL, convex RL, and risk-sensitive RL, see App. D. A diagram summarizing these connections is provided in App. G.2.
@@ 7 Conclusion
We introduced a theoretical policy-gradient framework for CPT-PO, paving the way for more efficient algorithms in practice. Looking ahead, an important direction is learning the utility and probability distortion functions directly from data to better align with human preferences; another is modeling collective human behavior for incentive design via CPT. By advancing behaviorally aligned RL, we aim to stimulate further research on more nuanced, human-aware decision making beyond traditional expected-utility approaches, bridging machine learning, behavioral economics, psychology, and real-world needs.
@@ Acknowledgements
Appendix A Broader Impact Statement
This paper contributes to the development of human-centered sequential decision-making models incorporating cognitive and psychological biases, which are essential for high-stakes, socially beneficial applications. Despite its importance, this perspective remains underexplored in RL, and our work helps fill this gap. We provide a theoretical framework for CPT-RL policy optimization and test it in simple applications for diverse domains such as energy, traffic control and finance, where improved decision-making can directly benefit society. Looking ahead, we envision that CPT-RL could be valuable in other human-centric fields such as legal and ethical decision-making, cybersecurity, and human-robot interaction, each carrying significant societal implications. Further practical research is required in this direction.
While our primary focus is theoretical and methodological, it is important to acknowledge potential risks, such as the challenge of accurately modeling human preferences or unintended biases introduced by learned CPT functions. Ensuring ethical, interpretable, and responsible deployment of CPT-RL in real-world applications is an essential direction for future research. By advancing behaviorally aligned RL, we hope this work contributes to more nuanced and human-aware decision-making beyond traditional utility-based approaches, helping to bridge the gap between machine learning, psychology, and real-world societal needs.
Appendix B Notation for Policy Classes
@@
Appendix C CPT-RL vs Risk-sensitive RL: An Illustration
  • Actions: The agent can move up, down, left, or right.
Results. See figures 5 and 6 and their captions in the next page.
Appendix D Extended Related Work Discussion
D.1 Risk-sensitive RL
@@
There is a rich literature around risk sensitive control and RL that we do not hope to give justice to here. We refer the reader to recent comprehensive surveys on the topic [Garcıa and Fernández, 2015, Prashanth et al., 2022] and the references therein. Let us briefly mention that there exist several approaches to risk sensitive RL. These include formulations such as constrained stochastic optimization to control the tolerance to perturbations and stochastic minmax optimization to model robustness with respect to worst case perturbations for instance. Another approach which is more relevant to our paper discussion consists in regularizing or modifying objective functions. Such modifications are based on considering different statistics of the return deviating from the standard expectation such as the variance or the conditional value at risk (e.g. Tamar et al. [2012], Chow and Ghavamzadeh [2014], Chow et al. [2018]) or even considering the entire distribution of the returns like distributional RL [Bellemare et al., 2023]. Another popular objective modification consists in maximizing an exponential criterion (e.g. Borkar [2002], Noorani et al. [2022]) to obtain robust policies w.r.t noise and perturbations of system parameters or variations in the environment. Noorani et al. [2022] designed a model-free REINFORCE algorithm and an actor-critic variant of the algorithm leveraging an (approximate) multiplicative Bellman equation induced by the exponential objective criterion. Moharrami et al. [2024] recently proposed and analyzed similar PG algorithms for the same exponential objective. Vijayan and LA [2023] introduced a PG algorithm for solving risk-sensitive RL for a class of smooth risk measures including some distortion risk measures and a mean-variance risk measure. Their approach is based on simultaneous perturbation stochastic approximation (SPSA) [Bhatnagar et al., 2013] using zeroth-order information to estimate gradients. Our CPT-PO problem covers several of the aforementioned objectives including smooth distortion risk measures and exponential utility as particular cases (see App. G for more details).
D.2 Convex RL/RL with General Utilities
@@
In the last few years, convex RL (a.k.a. RL with general utilities) [Hazan et al., 2019, Zhang et al., 2020, Zahavy et al., 2021, Geist et al., 2022] has emerged as a framework to unify several problems of interest such as pure exploration, imitation learning or experiment design. More precisely, this line of research is concerned with maximizing a given functional of the state(-action) occupancy measure w.r.t. a policy. To solve this problem, several policy gradient algorithms have been proposed in the literature [Zhang et al., 2021, Bai et al., 2022, Barakat et al., 2023, 2024]. Mutti et al. [2022b, a, 2023a] challenged the initial problem formulation and proposed a finite trial version of the problem which is closer to practical concerns as it consists in maximizing a functional of the empirical state(-action) distribution rather than its true asymptotic counterpart. The particular case of our CPT policy optimization problem without probability distortion (see EUT-PO below) coincides with a particular case of the single trial convex RL problem [Mutti et al., 2023b] in which the function of the empirical visitation measure is a linear functional of the reward function (see App. G.6 for details). However, our general problem is not a particular case of convex RL which does not account for probability distortions. Furthermore, our utility function is in general nonconvex in our setting (see example in Fig 8) and our policy gradient algorithm is model-free. More recently, De Santi et al. [2024] introduced a global RL problem formulation where rewards are globally defined over trajectories instead of locally over states and used submodular optimization tools to solve the resulting non-additive policy optimization problem. While global RL allows to account for trajectory-level global rewards, it does not take into consideration probability distortions. In addition, their investigation is restricted to the setting where the transition model is known whereas our PG algorithm is model-free.
D.3 Cumulative Prospect Theoretic RL
@@
Motivated by Prospect Theory and its sibling CPT [Kahneman and Tversky, 1979, Tversky and Kahneman, 1992, Barberis, 2013], L.A. et al. [2016] first proposed to combine CPT with RL to obtain a better model for human decision making. Following this first research effort, only few isolated works [Borkar and Chandak, 2021, Ramasubramanian et al., 2021, Ethayarajh et al., 2024] considered a similar CPT-RL setting. In particular, Borkar and Chandak [2021] proposed and analyzed a Q-learning algorithm for CPT policy optimization. Ramasubramanian et al. [2021] further developed value-based algorithms for CPT-RL by estimating the CPT value of an action in a given state via dynamic programming. More precisely, they were concerned with maximizing a sum of CPT value period costs which is amenable to dynamic programming. In contrast to their accumulated CPT-based cost (see their remark 1), our CPT policy optimization problem formulation is different: we maximize the CPT value of the return of a policy (see CPT-PO). In particular, this objective does not enjoy an additive structure and hence does not satisfy a Bellman equation. Moreover, their work relying on value-based methods is restricted to finite discrete state action spaces. Our PG algorithm is also suitable for continuous state action settings as we demonstrate in our experiments. More recently, Ethayarajh et al. [2024] incorporated CPT (without probability distortion) into RL from human feedback for fine-tuning large language models. CPT has also been recently exploited for multi-agent RL [Danis et al., 2023]. Our work is complementary to this line of research, especially to L.A. et al. [2016] and its extended version Jie et al. [2018] which are the most closely related work to ours. While their algorithm design makes use of simultaneous perturbation stochastic approximation (SPSA) [Spall, 1992] using only zeroth order information, we rather propose a PG algorithm exploiting first-order information thanks to our special problem structure involving the CPT value of a cumulative sum of rewards. See section 3 for further details regarding this comparison.
@@
We refer the reader to App. G.2 for a summarizing diagram illustrating the relationships between CPT-RL, convex RL and risk-sensitive RL.
Appendix E Applications of CPT: From Prior Work in Stateless Settings to CPT-RL
E.1 Energy: Renovation and Home Energy Management
E.2 Security: Building Evacuation
E.3 Urban Planning and Mobility
E.4 Finance
E.5 Example: Personalized Treatment for Pain Management
E.6 Further Applications of CPT-RL
  • Electricity management. Please see simulations in the main part (section 5 and App. K.7) in a simple example setting to illustrate our methodology.
  • Finance: portfolio optimization, risk management, behavioral asset pricing (e.g. influence of investor sentiment on price dynamics via e.g. over-weighting of low-probability events, including their preferences). For recent applications of CPT to finance, we refer the reader to a recent paper Luxenberg et al. [2024] using CPT for portfolio optimization (in a stateless static setting). We also applied our methodology to financial trading (see App. K.8).
Appendix F CPT-RL and Trajectory-Based Reward RL as Preference Learning Paradigms
  1. (a)
  2. (b)
  3. (c)
  4. (d)
Appendix G More About CPT Values and CPT Policy Optimization
G.1 More about Optimal Policies in CPT-RL
The need for stochastic policies. We start our discussion by pointing out a stark difference between optimal policies in standard MDPs and CPT-PO. While there exists an optimal deterministic stationary policy for MDPs (see e.g. Thm. 6.2.10 in Puterman [2014]), this is not the case in general for CPT-PO. Indeed, there does not always exist an optimal policy for CPT-PO in $\Pi_{NM}^{D}\,.$ In general, stochasticity of the policy is essential in solving our CPT-RL problem. To see this, consider an example built around a function $w^{+}$ that puts special emphasis on the $10\%$ of the best outcomes. In this case, the optimal policy needs to be randomized to take advantage of this and obtain the highest returns with some probability without suffering from bad outcomes by deterministically committing to this riskier strategy (see App. H.1 for a proof).
G.2 Positioning CPT-RL in the literature
Remark 8.
G.3 CPT value examples
Setting Utility function $^{+} 𝒘w^{-}$
CPT Any Any Any
EUT Any Identity function Identity function
Distortion risk measure Identity function Any $1-w^{+}(1-t)$

### v4

@@ 5 Numerical Simulations
Setting: We train a softmax policy using CPT-PG, vPG and ERS-PG. Vanilla PG optimizes for the standard expected return (identity for $u$ and $w$), CPT-PG using a KT model (as defined in section 2-2.1) with S-shaped utility using parameters ($\alpha=0.6$, $\lambda=2.5$) and an S-shaped probability function $w$ under-weighting the probability 1/2. ERS-PG is run using an exponential (concave) risk sensitive utility $u(x)=\eta^{-1}(1-\exp(-\eta x))$ with $\eta=0.5\,.$ The reference point is set to be $x_{0}=0$ in this experiment. We use the exact same parameters for both environments.
Results and interpretation: As shown in Fig. 1 (right), in the Gain Bandit vPG selects the risky arm (higher expected return), whereas both CPT-PG and ERS-PG choose the safe arm—reflecting human risk aversion over gains. In the Loss Bandit, only CPT-PG “flips” to the risky arm, capturing human risk-seeking over losses. Crucially, CPT-PG does this with the same S-shaped utility and probability-weighting parameters in both settings, whereas the concave ERS-PG objective remains risk-averse throughout. This exact “safe in gains, risky in losses” pattern—known as the reflection effect—is the hallmark of Prospect Theory (Kahneman and Tversky, 1979; Tversky and Kahneman, 1992) from which our simple example is inspired. Neither expected-utility nor a single-parameter risk-sensitive criterion can reproduce both behaviors simultaneously, but CPT-RL can, using a fixed, psychologically-grounded model. While risk-sensitive RL can, with appropriate parameter tuning, replicate some aspects of human decision-making, it often lacks the asymmetric treatment of gains and losses and the probabilistic distortion that characterize human behavior. Overall our main message is that CPT-RL offers more flexibility in modeling. Note also that CPT captures several risk measures as particular cases using discontinuous weights (see App. I.3I.4).
@@
(b) Robustness to state space size. We compare our PG algorithm with the zeroth-order (CPT-SPSA-G) of L.A. et al. (2016) on MDPs with increasing size $n\times n$. Fig. 2 shows that our CPT-PG algorithm scales better to larger grid sizes than CPT-SPSA-G as expected due to its use of (first-order) gradient information. See App. J.5 for details.
@@
(c) Application to finance. The goal is to train RL trading agents using our PG algorithm in the CPT-RL setting using a gym trading environment and data from the Bitcoin USD market. See App. J.8 for more details. We test several utility and probability weighting functions including a risk averse exponential of the form $x\mapsto\frac{1}{\beta}(1-\exp(-\beta x))$ with different values of $\beta$ as well as the KT (Kahneman and Tversky) function with different values of the reference point $x_{0}$ to illustrate its influence. In Figure 3, we make three observations. First, the reference point shifts the values of the achieved CPT returns: The smaller the reference point, the larger are the returns (Fig. 3, left). This is because only values larger than the reference point are perceived as positive returns. This illustrates how the subjective perception of the agent of the returns is taken into account by the model. Second, different values of $\beta$ lead to different trajectories overall which can translate to different levels of risk aversion. In particular, the curves do not match the identity utility case in the first episodes and show more or less risk taken towards optimizing the CPT returns. Third, the exponent $\alpha$ in the utility distorts the function and shifts the returns significantly (Fig. 3, right). Lower values of $\alpha$ lead to higher returns in this setting. This parameter $\alpha$ provides a degree of freedom to model the behavior of the agent as per their perception of the returns. Different values of $\alpha$ modify the curvature of the utility function (w.r.t. the reference point $x_{0}=0$ here) which is concave for gains and convex for losses.
@@
Remark 8.
(Reference point). The reference point is typically learned from the data in specific human-related applications by personalized tuning. In our experiment (see Fig. 3), we vary this reference point to show how the performance is immediately influenced by this parameter of the model. If the agent has different perceptions of gains and losses with a different reference/tolerance point, then policy optimization takes this into account.
Remark 9.
(History dependent policies and non-Markovianity). In our experiments, we incorporated partial history by expanding the input window of the networks to represent multiple past states. Typically, the input size was chosen to capture sufficient context without encoding the full trajectory length. This limited form of history dependence was sufficient in our settings, although incorporating more expressive temporal models could further enhance performance. The question of scaling to higher dimensional problems deserves further investigation. It is worth noting that the memory requirement is heavily correlated with the temporal structure of the data, especially in time series data like in finance. In our simple simulations, a small expanded input window was enough to obtain a descent performance. Our main goal was to show sensitivity of the KT model to different hyperparameters rather than optimizing performance.
Remark 10.
(Markovian vs non-Markovian policies). In App. J.4 (see Fig. 22), we provide a simple example where we show how a non-Markovian policy performs better than a Markovian one when running CPT-PG.
Additional simulations. We provide more simulations demonstrating the applicability of our CPT-PG algorithm in different settings, including illustrations of our theoretical results (App. J.3-J.4) and applications to a traffic control example over a grid (App. J.6), an electricity management example (App. J.7) and a control application (App. J.9).
@@ 6 Related Work
Prospect Theory and its sibling, CPT (Kahneman and Tversky, 1979; Tversky and Kahneman, 1992; Barberis, 2013), were first integrated with RL by L.A. et al. (2016). Since then, only a few studies have explored the CPT-RL framework (Borkar and Chandak, 2021; Ramasubramanian et al., 2021; Ethayarajh et al., 2024). Notably, Borkar and Chandak (2021) proposed a Q-learning algorithm for CPT-based policy optimization, while Ramasubramanian et al. (2021) developed value-based methods estimating CPT values using dynamic programming. Their approach optimizes a sum of CPT-transformed period costs, making it amenable to dynamic programming (see remark 1 therein). In contrast, our CPT formulation is different: we maximize the CPT value of the return of a policy CPT-PO. This objective lacks an additive structure, hence does not satisfy a Bellman equation, rendering dynamic programming approaches inapplicable. Additionally, prior value-based methods are limited to finite state-action spaces, whereas our PG algorithm is also suitable for continuous state action settings, as shown in our experiments. More recently, Ethayarajh et al. (2024) incorporated CPT (without probability distortion) for fine-tuning large language models with human feedback. Our work complements prior CPT-RL studies (L.A. et al., 2016; Jie et al., 2018) that rely on zeroth-order SPSA methods (Spall, 1992). Instead, we introduce a PG algorithm that leverages first-order information, exploiting the structure of CPT values applied to cumulative rewards (see Section 3 for further comparison). Unlike existing PG approaches in risk-sensitive RL, our method explicitly accounts for probability distortion and S-shaped utility transformations, key aspects of CPT. For the special case of DRMs, as previously discussed in more details, Vijayan and L.A (2024) proposed a policy gradient method for maximizing DRM objectives and provided non-asymptotic first-order stationary guarantees. Recently, Pachal et al. (2025) proposed a policy Newton algorithm for maximizing DRM objectives and established a a non-asymptotic bound that establishes the convergence of the algorithm to approximate second-order stationary policies. We focus on first-order PG algorithms without resorting to higher-order Hessian information and we consider the more general class of CPT objectives.
For a broader discussion on CPT-RL, convex RL, and risk-sensitive RL, see App. B. A diagram summarizing these connections is provided in App. I.2.
@@ 7 Conclusion
We developed a policy-gradient framework for CPT-based policy optimization in finite-horizon MDPs, including a policy gradient theorem, a Monte Carlo gradient estimator, and convergence guarantees for a first-order CPT-PG algorithm. Our simulations illustrate qualitative behaviors induced by CPT objectives and compare first-order updates to existing zeroth-order approaches. A natural direction for future work is to relax the assumption that the utility and probability-distortion functions are specified, e.g., by learning or calibrating them from data and studying the resulting statistical and optimization trade-offs. Another direction is to extend CPT-based objectives to multi-agent or social settings, where reference points and probability distortions may interact with strategic incentives.
Broader Impact Statement
This work develops optimization tools for sequential decision-making objectives inspired by behavioral economics, specifically Cumulative Prospect Theory (CPT), which models probability distortion and asymmetric valuation around a reference point. Such objectives are relevant in settings where decision-makers exhibit systematic deviations from expected-utility assumptions, including human-in-the-loop or preference-driven applications. Our primary contribution is theoretical and methodological: we provide a policy-gradient framework for CPT-RL and demonstrate it in simulation environments drawn from several application themes (e.g., finance, traffic control, and energy) as illustrative case studies.
Potential risks include mis-specification of CPT components (utility and distortion functions), reinforcement of undesirable biases if these components are learned from data, and misuse of behavioral models to manipulate users. Any deployment in real-world, human-facing systems would require careful calibration and validation, transparency about modeling assumptions, and appropriate safeguards to ensure responsible use.
@@ Acknowledgements
Appendix A Notation for Policy Classes
@@
Appendix B Extended Related Work Discussion
B.1 Risk-sensitive RL
@@
There is a rich literature around risk sensitive control and RL that we do not hope to give justice to here. We refer the reader to recent comprehensive surveys on the topic [Garcıa and Fernández, 2015, Prashanth et al., 2022] and the references therein. Let us briefly mention that there exist several approaches to risk sensitive RL. These include formulations such as constrained stochastic optimization to control the tolerance to perturbations and stochastic minmax optimization to model robustness with respect to worst case perturbations for instance. Another approach which is more relevant to our paper discussion consists in regularizing or modifying objective functions. Such modifications are based on considering different statistics of the return deviating from the standard expectation such as the variance or the conditional value at risk (e.g. Tamar et al. [2012], Chow and Ghavamzadeh [2014], Chow et al. [2018]) or even considering the entire distribution of the returns like distributional RL [Bellemare et al., 2023]. Another popular objective modification consists in maximizing an exponential criterion (e.g. Borkar [2002], Noorani et al. [2022]) to obtain robust policies w.r.t noise and perturbations of system parameters or variations in the environment. Noorani et al. [2022] designed a model-free REINFORCE algorithm and an actor-critic variant of the algorithm leveraging an (approximate) multiplicative Bellman equation induced by the exponential objective criterion. Moharrami et al. [2024] recently proposed and analyzed similar PG algorithms for the same exponential objective. Vijayan and L.A. [2023] introduced a PG algorithm for solving risk-sensitive RL for a class of smooth risk measures including some distortion risk measures and a mean-variance risk measure. Their approach is based on simultaneous perturbation stochastic approximation (SPSA) [Bhatnagar et al., 2013] using zeroth-order information to estimate gradients. Our CPT-PO problem covers several of the aforementioned objectives including smooth distortion risk measures and exponential utility as particular cases (see App. I for more details).
B.2 Convex RL/RL with General Utilities
@@
In the last few years, convex RL (a.k.a. RL with general utilities) [Hazan et al., 2019, Zhang et al., 2020, Zahavy et al., 2021, Geist et al., 2022] has emerged as a framework to unify several problems of interest such as pure exploration, imitation learning or experiment design. More precisely, this line of research is concerned with maximizing a given functional of the state(-action) occupancy measure w.r.t. a policy. To solve this problem, several policy gradient algorithms have been proposed in the literature [Zhang et al., 2021, Bai et al., 2022, Barakat et al., 2023, 2025]. Mutti et al. [2022b, a, 2023] challenged the initial problem formulation and proposed a finite trial version of the problem which is closer to practical concerns as it consists in maximizing a functional of the empirical state(-action) distribution rather than its true asymptotic counterpart. The particular case of our CPT policy optimization problem without probability distortion (see EUT-PO below) coincides with a particular case of the single trial convex RL problem [Mutti et al., 2023] in which the function of the empirical visitation measure is a linear functional of the reward function (see App. I.6 for details). However, our general problem is not a particular case of convex RL which does not account for probability distortions. Furthermore, our utility function is in general nonconvex in our setting (see example in Fig 11) and our policy gradient algorithm is not model-based in the sense that we do not estimate the state transition model. More recently, De Santi et al. [2024] introduced a global RL problem formulation where rewards are globally defined over trajectories instead of locally over states and used submodular optimization tools to solve the resulting non-additive policy optimization problem. While global RL allows to account for trajectory-level global rewards, it does not take into consideration probability distortions. In addition, their investigation is restricted to the setting where the transition model is known whereas our PG algorithm does not require to know or estimate the transition model.
B.3 Cumulative Prospect Theoretic RL
@@
Motivated by Prospect Theory and its sibling CPT [Kahneman and Tversky, 1979, Tversky and Kahneman, 1992, Barberis, 2013], L.A. et al. [2016] first proposed to combine CPT with RL to obtain a better model for human decision making. Following this first research effort, only few isolated works [Borkar and Chandak, 2021, Ramasubramanian et al., 2021, Ethayarajh et al., 2024] considered a similar CPT-RL setting. In particular, Borkar and Chandak [2021] proposed and analyzed a Q-learning algorithm for CPT policy optimization. Ramasubramanian et al. [2021] further developed value-based algorithms for CPT-RL by estimating the CPT value of an action in a given state via dynamic programming. More precisely, they were concerned with maximizing a sum of CPT value period costs which is amenable to dynamic programming. In contrast to their accumulated CPT-based cost (see their remark 1), our CPT policy optimization problem formulation is different: we maximize the CPT value of the return of a policy (see CPT-PO). In particular, this objective does not enjoy an additive structure and hence does not satisfy a Bellman equation. Moreover, their work relying on value-based methods is restricted to finite discrete state action spaces. Our PG algorithm is also suitable for continuous state action settings as we demonstrate in our experiments. More recently, Ethayarajh et al. [2024] incorporated CPT (without probability distortion) into RL from human feedback for fine-tuning large language models. CPT has also been recently exploited for multi-agent RL [Danis et al., 2023]. Our work is complementary to this line of research, especially to L.A. et al. [2016] and its extended version [Jie et al., 2018] which are the most closely related work to ours. While their algorithm design makes use of simultaneous perturbation stochastic approximation (SPSA) [Spall, 1992] using only zeroth order information, we rather propose a PG algorithm exploiting first-order information thanks to our special problem structure involving the CPT value of a cumulative sum of rewards. See section 3 for further details regarding this comparison.
@@
We refer the reader to App. I.2 for a summarizing diagram illustrating the relationships between CPT-RL, convex RL and risk-sensitive RL.
Appendix C Proofs for Section 2.3
C.1 The need for stochastic policies
To prove the result (i.e. the need for stochastic policies), we consider a simple MDP with only two states (an initial state and a terminal one) and two actions (A and B). See Fig. 14(a) below. We choose the identity as utility. Action A yields reward 1 with probability 1 and action B yields either 0 or $\frac{3}{2}$ with probability $\frac{1}{2}$ each. We further consider the following probability distortion function $w_{+}:[0,1]\to[0,1]$ defined for every $x\in[0,1]$ as follows: